Use the formula for finding a z-score to determine the missing value in the following table. Round your answer to two decimal places, if necessary.

z x μ σ
1.14 ? −23.40 3.28
Answer #1

z = \frac{X - \mu}{\sigma}

Given, z = 1.14

\mu = -23.40

\sigma = 3.28

X = -23.40 + 1.14*3.28 = -19.66